Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Lock acquisition: auxiliary lasers)
bersanetti, berni, casanueva, mantovani, masserot, pinto - 16:35 Monday 11 January 2021 (50409) Print this report
Lock of NARM and WARM with ALS, OLTF measurements

This morning we worked on the lock of the ALS system on both arms:

  • the new implementation of the dac_offset is not a constant value anymore but a ramp, which can be slowed down and then stopped in order to engage the lock in the correct working point (similar in concept to the lock of the OMC); this uses two relays ALS_[N,W]EB_DC[1,2]_sw, which need to be set with the same thresholds. Given the automatic management from the dac_offset ACL_RAMP_CH object, the relays do not need to be set manually to automatic mode (-2) as this is automatically handled by the dac_offset itself, which is the one we need to operate on;
  • the dac_offset object can be disabled or put in 'scan' or 'lock' mode (-1, 0 and 1 respectively), which is the command that engages the lock (alongside with setting the gain, as usual); there are buttons on the two VPM processes, or the following cm command has to be used (e.g. for the West arm); factor is the slowing down ratio of the ramp when the first relay goes to 1 (default is 4), while mode is the [-1,0,1] parameter described above. Both must be written as integers:
    • cm_send('WEB_ALS_fast','AcRampChTranSet','dac_offset',factor,mode)
  • with the new implementation we managed to lock both arms (with the IR locked) and we injected sensing noise in order to measure the OLTF of the two loops; results are in Figures 1 and 2 (North and West respectively); we have a UGF of about 3 kHz, and there is the same structure on both of them, around 5.7 kHz, although the one on the West arm is less visible; we will study this feature of the loops, but we do not have much phase, so we will see how to increase the UGF;
  • with the West arm, we managed to stay locked with ALS after unlocking the IR, with the DAC having to deal with a lot of low frequency, but keeping the lock (13:17:10 UTC + 180 s, Figure 3);
  • after this, we managed to acquire the lock directly with ALS without the IR beam locked (Figure 4);
  • during the shift, we found out that we have difficulties in working and locking both arms at the same time; to overcome this, all the processes relative to the West Arm are being moved from rtpc22 to rtpc24, so that both can be used simultaneously.
Images attached to this report
Comments to this report:
masserot - 16:49 Monday 11 January 2021 (50412) Print this report

The WEB ALS servers are now running on the rtpc24.

The ALS_{NEB,WEB}_Err_post_10KHz and the ALS_{NEB,WEB}_Corr_10KHz are now sent to the LSC rtpc.

Search Help
×

Warning

×